the width of a rectangle is 2 meters more than one third of the length . the perimeter of the rectangle is 44 meters. find the length and width of the rectangle.
thank you

Answer by anshu_a(16)

let lenth = x m, width = x/3+2 m ;
perimeter 2(x+x/3+2) = 44
x+x/3+2 = 22
4x/3 = 20
x = 20*3/4 = 15
lenth = 15
width = 15/3+2 = 7